FantasySadness:

Alex has 36 trophies in total. He has 2/3 trophies for track. He has 1/9 of the trophies for basketball. The remaining is for football. How many trophies does he have for football?? PLS HELP!

jhonyy9:

total 36 trophies \[\frac{ 2 }{ 3 }36 \] for track \[\frac{ 1 }{ 9 }36 \] for basketball R the remaining for football so using these above wrote we can write an equation \[\frac{ 2 }{ 3 }36 + \frac{ 1 }{ 9 }36 + R = 36 \] and now just solve this above wrote equation for R and in this way will get easy right sure the number of trophies what he have for football hope helped understandably easy